find the equation of line whose distance from origin is 2unit and passes through (2,2)

Grade:11

1 Answers

Aditya dhonde
7 years ago
We know that, y=mx+c here m=1 i.e slope of the line c=2 i.e. y-intercept therefore eq.n is y=x+2 or x-y+2=0